OpenStudy (anonymous):

please solve for: 3x squared - 9x + 3 when x = 1/2

OpenStudy (anonymous):

0

OpenStudy (anonymous):

owaitnvm

OpenStudy (anonymous):

just plug in 1/2 in place of x

OpenStudy (anonymous):

3/4 XD

OpenStudy (anonymous):

\[3 \times \frac{1}{4} - \frac{9}{2} +3\] \[\frac{-3}{4}\]